Suzaku, Courtesy JAXA

X-ray Imaging Spectrometer (XIS)
News

Suzaku News

For all the current news about the Suzaku X-ray satellite, visit the What's New page at the Suzaku GSFC Guest Observer Facility and the Suzaku Documents page at ISAS.

XIS News

January 5, 2010 - XIS0 Charge Leakage Anomaly Final Report

Final report on the XIS0 anomaly released as Suzaku Memo #2010-01 (PDF).

July 2, 2009 - XIS0 Charge Leakage Anomaly

From the NASA Guest Observer Facility (GOF):

On June 23, 02:01 UT, XIS0 suddenly showed an anomaly in Segment A. A large amount of charge produced in the imaging region filled several tens of columns in the segment. The XIS frame consists of 4 segments, A through D, and only a half of the segment, i.e. 1/8 of the whole imaging region, is affected. The unusable columns are 70-150 in the ACTX coordinate, and are thus at the edge of the field of view. The phenomenon looks similar to one that occurred in XIS2 in November, 2006, but, fortunately, the leakage of charge is much smaller this time. We are trying to identify the cause of this anomaly, and to reduce the charge leakage. Sensors XIS1 and XIS3 are working fine.

Meanwhile, we will continue the Suzaku observations as planned, because we believe nominal XIS operations are unlikely to cause any further degradation. We will mask an region 80 columns wide surrounding the affected area to prevent spurious events from filling the telemetry. Although the planned observations are being conducted, usage of the burst/window options is not supported for XIS0 until the XIS team confirms the operation of XIS0 in these modes. This confirmation will be done soon, before the start of the next observations requiring the burst/window options.

Although the anomaly is restricted in Segment A, it could affect the calibration of other segments slightly. The XIS team will perform additional calibration observations, if necessary, and will update the calibration data appropriately.

Kazuhisa Mitsuda
Suzaku Project Manger

December 17, 2007 - XIS0 Contamination Increase

From the XIS Team:

The XIS team has found that the contamination of XIS0 is increasing more rapidly than expected.

September 20, 2007 - XIS Dark Frame Processing Error

From the XIS Team:

It was found that the dark frame data used for the on-board processing became erroneous in some occasions. The error was found some time between April 29 and October 10, 2006, and only the event data in the region of RAWY < 119 were affected. Thus the impact on point source observations was limited, but extended sources and background observations were affected. We describe below the nature of this error and how it affects the X-ray data.

November 23, 2006 - XIS2 Anomaly and Status

From the GOF:

On November 9, about 2/3 of imaging area of XIS2, one of the X-ray CCD cameras, became suddenly unusable. All other sensors, XIS0, 1, 3, are fine. Since there was a concern that the start of spaced-row charge injection (SCI) might have triggered the problem, we temporally stopped the SCI just after the trouble. However, after extensive analysis and in-orbit and ground tests, we concluded the SCI is not related to the trouble. We will soon resume the SCI step by step.

Our attempts to recover XIS2 have so far proven unsuccessful, but we will continue our efforts to identify the cause of the problem. In the mean time, we will increase the exposure times of Priority A targets of cycle-1 GO observations by 25 %, and use them as the observation-completion criteria. However we will not increase those of Priority B and C targets, thus the observation-completion criteria for Priorities B and C remain the same.

We ask cycle-2 proposers to perform proposal simulations assuming four XIS sensors. We will consider adjusting observation times according to evaluation by the selection boards.

best regards,
Kazuhisa Mitsuda
Suzaku Project Manger

December 20, 2005 - Low Energy Efficiency Degradation

From the GOF:

The good low energy response of the back illuminated (BI) CCD, i.e. the high detection efficiency and the gaussian-like good energy spread function, is an important unique feature of Suzaku. However, we recently found evidence that the effective area below 1 keV of both back- and front-illuminated sensors is decreasing with time (see figures). On the average at 0.6 keV the effective area is about 50% that just after launch. We are working hard to understand the cause of decrease and to find possible recovery methods. The present best estimate of the e-folding time constant is 160 days at the OVII line (570 eV), but it can be shorter at lower energies and is longer at higher energies. However, at present we cannot predict the future long-term behavior of efficiency well. Because of this we will put the priority on the low energy observations as soon as possible during AO1. And we will provide the response functions (rmf files) of the XIS as of October 26, 2005 on the Suzaku web page soon. Please note in your proposal which response function is used to estimate the observation times.

Please note that the energy resolution has not changed.

Best wishes,
Kazuhisa Mitsuda (ISAS/JAXA)
Hideyo Kunieda (ISAS/JAXA and Nagoya University)
Nick White (NASA/GSFC)


August 13, 2005

The remaining three XIS doors were opened with no problems. The counting rates from E0102-72 are about as expected, and the sensors seem reasonably well-aligned with one another. The performance of the door's paraffin actuator was nominal in each case.

August 12, 2005 - Suzaku First Light

The first XIS door (XIS3) was opened today, and X-rays from the supernova remnant E0102-72 were observed. It is the first light for the Suzaku satellite! Count rates, spectral resolution and PSF are close to the expected values, albeit from a single orbit worth of data.

August 1, 2005

XIS TEC hot-side temperatures have been reduced to -43C or so; these temperatures are controlled by means of heaters on the XIS heatpipes. As a result, TEC voltage is down to 4.8V to 5V. Various XIS timing modes (including parallel-sum, window- and burst-modes) are being tested.

July 28, 2005

Last night we downlinked 5x5 event mode data that had been accumulating during remote (non-contact) orbits. A quick look at the data from two of the four sensors (XIS0=w1.3c6 FI, and XIS1=w1.8c2 BI). We don't have many counts because the calibration source on the door produces only a few counts/CCD/s, but the resolution at 5.9 keV is in the range 130-140 eV FWHM for all 8 quadrants checked so far, as expected (see the Calibration section).

July 27, 2005

All seems well with the XIS. The TECs were turned on last night and the CCDs cooled to -60C. Doors are closed and cal sources are illuminating the CCDs. During this evening's passes:

  • CCD temperature was lowered in stages to (nominally) -90C. Control was obtained on all 4 at -89 < T < -90 with the TEC hot side at -36C to -38C and 5.6 to 5.7 V TEC voltage.
  • We looked at some frame mode data at -60, -80 and -90 C. Noise on the quads checked so far (including BI = XIS1) is in the range 2-2.2 ADU (roughly 1 ADU = 1 electron) both in overclock and imaging areas. General appearance was as expected, with larger particle images in FI than BI (see the Calibration section for sample images).
  • DE/DP successfully computed a dark/bias frame
  • We acquired 5x5 event mode data; as expected, raw (unrejected) event rate is higher in FI than BI.


MKI     
MIT
Last updated: Fri Jan 29 15:13:42 EST 2010
email: milleric@mit.edu